What is snell law and write an expression.


1
Expert's answer
2021-01-04T14:39:37-0500

Snell's law determines the direction of propagation of a ray of light that falls on the interface between two different isotropic media


"\frac{\sin\alpha_2}{\sin\alpha_1}=\frac{v_2}{v_1}=\frac{n_1}{n_2}"